# Multiview Random Vector Functional Link Network for Predicting DNA-Binding Proteins

### v1

@@ A.2.2 Performance with different number of training samples on AwA dataset
Figure 11: Performance with different number of training samples of the proposed MvRVFL model on AwA dataset.
A.3 Effect of hyperparameters $C_{1}$ and $\sigma$
A.3.1 Sensitivity of $\theta$ and $\rho$ combinations
@@
We explore the sensitivity of the MvRVFL model to various values of parameters $\theta$ and $\rho$. The parameter $\theta$ regulates the gap between two views, while $\rho$ is linked to the coupling term $\xi_{1}^{t}\xi_{2}$. We vary the values of $\theta$ and $\rho$ from $10^{-5}$ to $10^{5}$ and record the corresponding Acc. results. With other parameters held constant at their optimal settings, Fig. 12 illustrates how the performance of MvRVFL changes as the values of parameters $\theta$ and $\rho$ are varied.
@@
(a) 1000
(b) 10000
(c) 103000
(d) 143000
Figure 12: Performance variation with different values of parameters $\theta$ and $\rho$ of the proposed MvRVFL model on Corel5k dataset.
A.4 Classification accuracy tables of the proposed MvRVFL models along with baseline models on UCI, KEEL, AwA and Corel5k datasets
@@
Table 7: Performance comparison of the proposed MvRVFL along with the baseline models based on classification Acc. for UCI and KEEL datasets.
Dataset SVM2K MvTSVM ELM-Vw-$A$ ELM-Vw-$B$ RVFL-Vw-$A$ RVFL-Vw-$B$ MVLDM MvRVFL-1 MvRVFL-2
(Farquhar et al., 2005) (Xie & Sun, 2015) (Huang et al., 2006) (Huang et al., 2006) (Pao et al., 1994) (Pao et al., 1994) (Hu et al., 2024)
$(C_{1})$ $(C_{1},C_{2},D)$ $(C,N)$ $(C,N)$ $(C,N)$ $(C,N)$ $(C_{1},\nu_{1},\nu_{2},\sigma)$ $(C_{1},C_{2},\rho,N)$ $(C_{1},C_{2},\rho,N)$
aus $87.02$ $71.15$ $85.98$ $86.06$ $85.98$ $85.98$ $71.98$ $87.5$ $87.98$
$(0.001)$ $(0.00001,0.00001,0.00001)$ $(0.01,123)$ $(0.001,163)$ $(0.001,163)$ $(10,23)$ $(0.001,0.01,0.1,0.25)$ $(1,0.0001,0.00001,63)$ $(100,100000,1,43)$
bank $80.74$ $71.86$ $85.54$ $85.39$ $85.83$ $85.61$ $73.67$ $90.05$ $89.24$
$(0.001)$ $(0.00001,0.00001,0.00001)$ $(0.00001,3)$ $(10000,23)$ $(1000,163)$ $(10,123)$ $(0.1,0.001,0.001,4)$ $(100,0.1,0.0001,203)$ $(100,1000,0.0001,83)$
breast_cancer $62.45$ $55.58$ $69.77$ $65.12$ $67.44$ $66.28$ $70$ $70.93$ $75.58$
$(0.001)$ $(0.00001,0.00001,0.00001)$ $(1,83)$ $(0.01,43)$ $(0.0001,3)$ $(1,83)$ $(0.0001,0.1,0.001,4)$ $(0.1,0.0001,0.00001,43)$ $(0.0001,0.01,0.0001,43)$
breast_cancer_wisc $90.04$ $81.43$ $92.1$ $92.1$ $92.14$ $95.1$ $75$ $98.57$ $96.19$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,203)$ $(0.01,63)$ $(0.00001,23)$ $(0.0001,203)$ $(0.0001,0.0001,0.00001,0.25)$ $(0.00001,0.1,0.00001,163)$ $(100,0.001,10,23)$
breast_cancer_wisc_diag $95.49$ $88.6$ $92.42$ $97.08$ $95.83$ $96.49$ $93.15$ $98.25$ $97.08$
$(0.001)$ $(0.00001,0.00001,0.00001)$ $(0.01,163)$ $(0.001,163)$ $(0.001,163)$ $(100,3)$ $(0.001,0.0001,1,4)$ $(0.001,0.0001,0.0001,123)$ $(0.1,10,0.0001,83)$
breast_cancer_wisc_prog $58.33$ $58.33$ $68.33$ $68.33$ $73.33$ $68.33$ $71.17$ $73.33$ $73.33$
$(0.001)$ $(0.00001,0.00001,0.00001)$ $(0.01,203)$ $(0.01,83)$ $(100000,3)$ $(0.01,203)$ $(0.0001,0.0001,0.1,2)$ $(1,10,0.00001,23)$ $(10000,100,100,3)$
brwisconsin $97.56$ $61.95$ $97.07$ $96.1$ $97.07$ $96.59$ $95.59$ $96.1$ $97.56$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,143)$ $(0.001,63)$ $(0.001,63)$ $(100000,23)$ $(0.001,0.00001,0.1,4)$ $(0.00001,1,0.001,163)$ $(0.0001,1,0.00001,203)$
bupa or liver-disorders $54.8$ $42.31$ $63.46$ $65.38$ $63.46$ $66.35$ $55.34$ $70.19$ $64.42$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(1000,23)$ $(0.1,163)$ $(0.1,163)$ $(0.1,163)$ $(0.001,0.001,0.001,4)$ $(10,0.01,0.001,23)$ $(0.1,1,0.0001,83)$
checkerboard_Data $87.02$ $43.75$ $86.98$ $86.06$ $86.98$ $86.98$ $84.06$ $87.5$ $87.98$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.01,123)$ $(0.001,163)$ $(0.001,163)$ $(10,23)$ $(0.001,0.001,0.001,1)$ $(1,0.0001,0.00001,63)$ $(100,100000,1,43)$
chess_krvkp $80.45$ $82.35$ $95.62$ $93.33$ $95.62$ $94.68$ $97.7$ $96.77$ $94.58$
$(0.001)$ $(0.00001,0.00001,0.00001)$ $(100,203)$ $(100000,203)$ $(10000,203)$ $(100,203)$ $(0.1,0.1,0.1,4)$ $(0.1,0.0001,0.1,203)$ $(100000,0.00001,10,203)$
cleve $80$ $75.56$ $80$ $85.56$ $81.11$ $81.11$ $84.27$ $84.44$ $81.11$
$(0.1)$ $(0.00001,0.00001,0.00001)$ $(0.001,103)$ $(0.001,123)$ $(0.01,3)$ $(0.01,23)$ $(0.01,0.0.1,0.1,1)$ $(0.00001,1,0.00001,203)$ $(0.0001,0.1,0.00001,123)$
cmc $64.25$ $55.88$ $69.91$ $70.14$ $68.1$ $71.72$ $74.38$ $72.17$ $70.36$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.01,143)$ $(0.01,183)$ $(100,63)$ $(1000,43)$ $(0.0001,0.01,0.001,4)$ $(1000,1,0.001,23)$ $(10000,0.00001,0.0001,23)$
conn_bench_sonar_mines_rocks $80.95$ $46.03$ $80.54$ $74.6$ $88.54$ $73.02$ $75.81$ $80.95$ $66.67$
$(1000)$ $(0.00001,0.00001,0.00001)$ $(10,3)$ $(0.01,203)$ $(0.01,203)$ $(0.1,183)$ $(0.0001,0.00001,10,0.5)$ $(10,1000,0.1,183)$ $(10,1000,10,163)$
cylinder_bands $68.18$ $60.39$ $76.62$ $74.03$ $72.08$ $74.68$ $71.9$ $74.68$ $78.57$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.001,203)$ $(100,143)$ $(0.001,183)$ $(1,23)$ $(0.0001,0.1,0.001,4)$ $(100,0.01,0.0001,43)$ $(10,1000,0.01,43)$
fertility $75.25$ $75$ $90$ $90$ $90$ $90$ $86.67$ $93.33$ $86.67$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.01,83)$ $(0.01,63)$ $(0.01,63)$ $(0.01,83)$ $(0.001,0.001,100,2)$ $(0.01,1,0.01,183)$ $(0.0001,10000,1,63)$
hepatitis $80.85$ $78.72$ $78.85$ $80.85$ $80.85$ $78.72$ $78.26$ $85.11$ $78.72$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(1,23)$ $(0.01,183)$ $(0.1,123)$ $(10,3)$ $(0.001,0.01,0.001,4)$ $(0.01,100,0.00001,203)$ $(0.01,100,0.00001,203)$
hill_valley $60.98$ $53.3$ $68.05$ $51.92$ $68.78$ $51.92$ $56.2$ $68.96$ $75.89$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(10000,103)$ $(1,103)$ $(10,103)$ $(10000,103)$ $(0.0001,0.0001,0.1,1)$ $(10,0.1,0.0001,123)$ $(100,1,100,143)$
mammographic $80.27$ $77.06$ $80.28$ $82.01$ $80.28$ $81.66$ $83.33$ $84.08$ $83.74$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(1,83)$ $(100000,63)$ $(100000,23)$ $(100,23)$ $(0.0001,0.001,1,4)$ $(1000,1,0.00001,23)$ $(1000,0.00001,0.0001,23)$
monks_3 $80.24$ $76.11$ $95.21$ $95.41$ $95.21$ $95.81$ $96.39$ $96.41$ $96.41$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.1,143)$ $(0.1,163)$ $(0.1,163)$ $(0.1,123)$ $(0.01,0.001,10,4)$ $(0.1,0.00001,0.001,163)$ $(0.1,1,0.00001,183)$
new-thyroid1 $78.46$ $82.31$ $98.46$ $96.92$ $98.46$ $96.92$ $95.31$ $100$ $100$
$(0.1)$ $(0.00001,0.00001,0.00001)$ $(10,103)$ $(1,43)$ $(1,23)$ $(10,103)$ $(0.0001,1000,0.1,0.25)$ $(0.001,0.0001,0.00001,183)$ $(1,0.001,0.01,123)$
oocytes_merluccius_nucleus_4d $74.27$ $64.82$ $82.41$ $81.11$ $83.71$ $80.78$ $75.16$ $82.08$ $81.76$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.1,83)$ $(1,143)$ $(1,143)$ $(0.1,83)$ $(0.1,100,0.1,4)$ $(10,100,0.1,143)$ $(1,10000,0.001,203)$
oocytes_trisopterus_nucleus_2f $78.83$ $58.39$ $86.13$ $78.83$ $85.04$ $78.83$ $82.05$ $85.04$ $86.5$
$(0.0001)$ $(0.00001,0.00001,0.00001)$ $(0.01,103)$ $(0.1,143)$ $(100000,43)$ $(0.01,103)$ $(0.1,0.001,1,4)$ $(100000,0.0001,100,83)$ $(0.1,100,0.1,203)$
parkinsons $71.19$ $71.19$ $81.36$ $77.97$ $89.83$ $77.97$ $93.1$ $89.83$ $93.22$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.1,163)$ $(1,63)$ $(0.1,203)$ $(0.1,163)$ $(0.0001,0.00001,1,2)$ $(1000,100000,0.1,163)$ $(1,1000,0.01,203)$
pima $76.19$ $73.33$ $74.89$ $74.03$ $74.46$ $74.03$ $69.13$ $76.19$ $74.89$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.1,63)$ $(0.01,143)$ $(0.01,143)$ $(0.1,63)$ $(0.0001,10,10,4)$ $(0.001,1,0.01,3)$ $(10000,0.01,0.001,3)$
pittsburg_bridges_T_OR_D $70.85$ $65.85$ $89.32$ $73.55$ $80.65$ $80.65$ $90$ $90.32$ $74.19$
$(100)$ $(0.00001,0.00001,0.00001)$ $(0.01,83)$ $(0.01,43)$ $(1000,3)$ $(0.01,3)$ $(0.1,0.00001,1,4)$ $(0.00001,1,0.01,163)$ $(10000,10,0.0001,3)$
planning $65.85$ $63.64$ $76.36$ $76.36$ $76.36$ $76.36$ $68.52$ $76.36$ $63.64$
$(10)$ $(0.00001,0.00001,0.00001)$ $(0.01,23)$ $(0.00001,3)$ $(0.00001,23)$ $(0.01,23)$ $(0.00001,0.001,1,1)$ $(0.01,0.00001,0.0001,23)$ $(0.0001,0.01,100000,43)$
ripley $89.07$ $80.67$ $86.13$ $87.53$ $87.53$ $87.53$ $89.07$ $89.87$ $89.6$
$(100000)$ $(0.00001,0.00001,0.00001)$ $(10000,103)$ $(100,203)$ $(10000,123)$ $(10000,103)$ $(0.001,0.001,0.001,4)$ $(1,0.00001,0.0001,123)$ $(10000,0.01,10,83)$
Average Acc. $76.65$ $67.24$ $82.66$ $80.95$ $83.14$ $81.26$ $79.9$ $85.15$ $83.18$
Average Rank $6.35$ $8.65$ $4.70$ $5.26$ $4.22$ $5.09$ $5.48$ $2.04$ $3.20$

### v2

@@ Appendix B Random Vector Functional Link (RVFL) Network
Figure S.1: Effect of parameter $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ on the performance of the proposed MvRVFL model on AwA dataset.
S.I.2 Performance with different number of training samples on AwA dataset
We assess how the proposed MvRVFL model’s performance varies with different numbers of training samples. Fig. S.2 shows how the Acc. changes as the number of training samples ranges from $86$ to $336$. The x-axis displays the number of training samples, and the y-axis shows the corresponding Acc. values. It is observed that the Acc. value generally increases with the rise in the number of training samples. This is because an increase in training samples provides more data for the model to learn from, leading to improved accuracy in the classification results.
Figure S.2: Performance with different number of training samples of the proposed MvRVFL model on AwA dataset.
S.I.3 Sensitivity of $\theta$ and $\rho$ combinations
@@
We explore the sensitivity of the MvRVFL model to various values of parameters $\theta$ and $\rho$. The parameter $\theta$ regulates the gap between two views, while $\rho$ is linked to the coupling term $\xi_{1}^{t}\xi_{2}$. We vary the values of $\theta$ and $\rho$ from $10^{-5}$ to $10^{5}$ and record the corresponding Acc. results. With other parameters held constant at their optimal settings, Fig. S.3 illustrates how the performance of MvRVFL changes as the values of parameters $\theta$ and $\rho$ are varied.
@@
Figure S.3: Performance variation with different values of parameters $\theta$ and $\rho$ of the proposed MvRVFL model on Corel5k dataset.
Appendix S.II Classification accuracy tables of the proposed MvRVFL models along with baseline models on UCI, KEEL, AwA and Corel5k datasets
@@
Table S.I: Performance comparison of the proposed MvRVFL along with the baseline models based on classification Acc. for UCI and KEEL datasets.
Dataset SVM2K [18] MvTSVM [70] ELM-Vw-AA [31] ELM-Vw-$B$ [31] RVFL-Vw-AA [46] RVFL-Vw-$B$ [46] MVLDM [29] CDC [34] MCMC [5] MvRVFL-1 MvRVFL-2
$(C_{1})$ $(C_{1},C_{2},D)$ $(C,N)$ $(C,N)$ $(C,N)$ $(C,N)$ $(C_{1},\nu_{1},\nu_{2},\sigma)$ $(\alpha,\beta,m)$ $(\lambda_{1},\lambda_{2},\lambda_{3},\zeta)$ $(C_{1},C_{2},\rho,N)$ $(C_{1},C_{2},\rho,N)$
aus $87.02$ $71.15$ $85.98$ $86.06$ $85.98$ $85.98$ $71.98$ $73.62$ $85.65$ $87.5$ $87.98$
$(0.001)$ $(0.00001,0.00001,0.00001)$ $(0.01,123)$ $(0.001,163)$ $(0.001,163)$ $(10,23)$ $(0.25,0.25,0.25,0.25)$ $(0.0001,10,100)$ $(0.01,1,0.0001,0.001)$ $(1,0.0001,0.00001,63)$ $(100,100000,1,43)$
bank $80.74$ $71.86$ $85.54$ $85.39$ $85.83$ $85.61$ $73.67$ $87.19$ $88.53$ $90.05$ $89.24$
$(0.001)$ $(0.00001,0.00001,0.00001)$ $(0.00001,3)$ $(10000,23)$ $(1000,163)$ $(10,123)$ $(0.25,0.25,0.25,4)$ $(1,10000,30)$ $(2.5,0.1,0.00001,0.0001)$ $(100,0.1,0.0001,203)$ $(100,1000,0.0001,83)$
breast_cancer $62.45$ $55.58$ $69.77$ $65.12$ $67.44$ $66.28$ $70$ $73.15$ $71.4$ $70.93$ $75.58$
$(0.001)$ $(0.00001,0.00001,0.00001)$ $(1,83)$ $(0.01,43)$ $(0.0001,3)$ $(1,83)$ $(0.25,0.5,0.5,4)$ $(10,1000,50)$ $(5,1,0.000001,0.00001)$ $(0.1,0.0001,0.00001,43)$ $(0.0001,0.01,0.0001,43)$
breast_cancer_wisc $90.04$ $81.43$ $92.1$ $92.1$ $92.14$ $95.1$ $75$ $93.65$ $95.22$ $98.57$ $96.19$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,203)$ $(0.01,63)$ $(0.00001,23)$ $(0.0001,203)$ $(0.5,0.25,0.25,0.25)$ $(1000,10,10)$ $(0.1,0.01,0.001,0.001)$ $(0.00001,0.1,0.00001,163)$ $(100,0.001,10,23)$
breast_cancer_wisc_diag $95.49$ $88.6$ $92.42$ $97.08$ $95.83$ $96.49$ $93.15$ $95.82$ $94.44$ $98.25$ $97.08$
$(0.001)$ $(0.00001,0.00001,0.00001)$ $(0.01,163)$ $(0.001,163)$ $(0.001,163)$ $(100,3)$ $(0.25,0.25,1,4)$ $(10000,1,100)$ $(1,2.5,0.0001,0.0001)$ $(0.001,0.0001,0.0001,123)$ $(0.1,10,0.0001,83)$
breast_cancer_wisc_prog $58.33$ $58.33$ $68.33$ $68.33$ $73.33$ $68.33$ $71.17$ $70.63$ $72.09$ $73.33$ $73.33$
$(0.001)$ $(0.00001,0.00001,0.00001)$ $(0.01,203)$ $(0.01,83)$ $(100000,3)$ $(0.01,203)$ $(0.25,0.25,4,2)$ $(0.0001,0.0001,70)$ $(10,5,0.000001,0.000001)$ $(0.00001,1000,0.01,23)$ $(10000,100,100,3)$
brwisconsin $97.56$ $61.95$ $97.07$ $96.1$ $97.07$ $96.59$ $95.59$ $95.34$ $92.74$ $96.1$ $97.56$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.0001,143)$ $(0.001,63)$ $(0.001,63)$ $(100000,23)$ $(0.25,0.25,2,4)$ $(1,10,100)$ $(0.01,10,0.00001,0.00001)$ $(0.00001,1,0.001,163)$ $(0.0001,1,0.00001,203)$
bupa or liver-disorders $54.8$ $42.31$ $63.46$ $65.38$ $63.46$ $66.35$ $55.34$ $65.07$ $66.14$ $70.19$ $64.42$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(1000,23)$ $(0.1,163)$ $(0.1,163)$ $(0.1,163)$ $(0.25,0.25,0.25,4)$ $(10,0.0001,30)$ $(2.5,2.5,0.0001,0.001)$ $(10,0.01,0.001,23)$ $(0.1,1,0.0001,83)$
checkerboard_Data $87.02$ $43.75$ $86.98$ $86.06$ $86.98$ $86.98$ $84.06$ $82.83$ $85.03$ $87.5$ $87.98$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.01,123)$ $(0.001,163)$ $(0.001,163)$ $(10,23)$ $(0.25,0.25,0.25,1)$ $(1000,10000,50)$ $(5,0.1,0.001,0.0001)$ $(1,0.0001,0.00001,63)$ $(100,100000,1,43)$
chess_krvkp $80.45$ $82.35$ $95.62$ $93.33$ $95.62$ $94.68$ $97.7$ $90.46$ $94.85$ $96.77$ $94.58$
$(0.001)$ $(0.00001,0.00001,0.00001)$ $(100,203)$ $(100000,203)$ $(10000,203)$ $(100,203)$ $(0.25,0.25,0.25,4)$ $(10000,10,10)$ $(0.1,5,0.0001,0.000001)$ $(0.1,0.0001,0.1,203)$ $(100000,0.00001,10,203)$
cleve $80$ $75.56$ $80$ $85.56$ $81.11$ $81.11$ $84.27$ $82.73$ $80.59$ $84.44$ $81.11$
$(0.1)$ $(0.00001,0.00001,0.00001)$ $(0.001,103)$ $(0.001,123)$ $(0.01,3)$ $(0.01,23)$ $(0.25,0.25,0.25,1)$ $(0.0001,1000,100)$ $(1,0.1,0.000001,0.00001)$ $(0.00001,1,0.00001,203)$ $(0.0001,0.1,0.00001,123)$
cmc $64.25$ $55.88$ $69.91$ $70.14$ $68.1$ $71.72$ $74.38$ $69.87$ $67.91$ $72.17$ $70.36$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.01,143)$ $(0.01,183)$ $(100,63)$ $(1000,43)$ $(0.25,0.25,0.25,4)$ $(1,0.0001,50)$ $(10,1,0.00001,0.001)$ $(1000,1,0.001,23)$ $(10000,0.00001,0.0001,23)$
conn_bench_sonar_mines_rocks $80.95$ $46.03$ $80.54$ $74.6$ $88.54$ $73.02$ $75.81$ $78.84$ $75.77$ $80.95$ $66.67$
$(1000)$ $(0.00001,0.00001,0.00001)$ $(10,3)$ $(0.01,203)$ $(0.01,203)$ $(0.1,183)$ $(0.25,0.25,4,0.5)$ $(10,10,70)$ $(0.01,0.01,0.000001,0.0001)$ $(10,1000,0.1,183)$ $(10,1000,10,163)$
cylinder_bands $68.18$ $60.39$ $76.62$ $74.03$ $72.08$ $74.68$ $71.9$ $72.45$ $75.18$ $74.68$ $78.57$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.001,203)$ $(100,143)$ $(0.001,183)$ $(1,23)$ $(0.25,4,0.25,4)$ $(1000,0.0001,30)$ $(2.5,10,0.000001,0.00001)$ $(100,0.01,0.0001,43)$ $(10,1000,0.01,43)$
fertility $75.25$ $75$ $90$ $90$ $90$ $90$ $86.67$ $86$ $92$ $93.33$ $86.67$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.01,83)$ $(0.01,63)$ $(0.01,63)$ $(0.01,83)$ $(0.25,0.25,2,2)$ $(10000,1000,50)$ $(5,2.5,0.00001,0.0001)$ $(0.01,1,0.01,183)$ $(0.0001,10000,1,63)$
hepatitis $80.85$ $78.72$ $78.85$ $80.85$ $80.85$ $78.72$ $78.26$ $75.82$ $74.19$ $85.11$ $78.72$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(1,23)$ $(0.01,183)$ $(0.1,123)$ $(10,3)$ $(0.25,0.25,0.25,4)$ $(0.0001,1,10)$ $(0.1,1,0.0001,0.001)$ $(0.01,100,0.00001,203)$ $(0.01,100,0.00001,203)$
hill_valley $60.98$ $53.3$ $68.05$ $51.92$ $68.78$ $51.92$ $56.2$ $65.33$ $67.82$ $68.96$ $75.89$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(10000,103)$ $(1,103)$ $(10,103)$ $(10000,103)$ $(0.25,0.25,0.5,1)$ $(1,1000,100)$ $(1,0.01,0.001,0.000001)$ $(10,0.1,0.0001,123)$ $(100,1,100,143)$
mammographic $80.27$ $77.06$ $80.28$ $82.01$ $80.28$ $81.66$ $83.33$ $81.61$ $83.26$ $84.08$ $83.74$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(1,83)$ $(100000,63)$ $(100000,23)$ $(100,23)$ $(0.25,0.25,1,4)$ $(10,1,100)$ $(10,0.1,0.000001,0.0001)$ $(1000,1,0.00001,23)$ $(1000,0.00001,0.0001,23)$
monks_3 $80.24$ $76.11$ $95.21$ $95.41$ $95.21$ $95.81$ $96.39$ $95.21$ $92.18$ $96.41$ $96.41$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.1,143)$ $(0.1,163)$ $(0.1,163)$ $(0.1,123)$ $(0.25,0.25,2,4)$ $(1000,1,70)$ $(0.01,0.1,0.00001,0.001)$ $(0.1,0.00001,0.001,163)$ $(0.1,1,0.00001,183)$
new-thyroid1 $78.46$ $82.31$ $98.46$ $96.92$ $98.46$ $96.92$ $95.31$ $98.16$ $98.67$ $100$ $100$
$(0.1)$ $(0.00001,0.00001,0.00001)$ $(10,103)$ $(1,43)$ $(1,23)$ $(10,103)$ $(0.25,2,0.5,0.25)$ $(10000,0.0001,10)$ $(2.5,0.01,0.0001,0.000001)$ $(0.001,0.0001,0.00001,183)$ $(1,0.001,0.01,123)$
oocytes_merluccius_nucleus_4d $74.27$ $64.82$ $82.41$ $81.11$ $83.71$ $80.78$ $75.16$ $80.54$ $81.38$ $82.08$ $81.76$
$(0.00001)$ $(0.00001,0.00001,0.00001)$ $(0.1,83)$ $(1,143)$ $(1,143)$ $(0.1,83)$ $(0.5,4,0.5,4)$ $(0.0001,10000,30)$ $(5,5,0.000001,0.00001)$ $(10,100,0.1,143)$ $(1,10000,0.001,203)$
oocytes_trisopterus_nucleus_2f $78.83$ $58.39$ $86.13$ $78.83$ $85.04$ $78.83$ $82.05$ $82.41$ $83.88$ $85.04$ $86.5$
$(0.0001)$ $(0.00001,0.00001,0.00001)$ $(0.01,103)$ $(0.1,143)$ $(100000,43)$ $(0.01,103)$ $(0.25,0.25,1,4)$ $(1,1,10)$ $(0.1,2.5,0.001,0.0001)$ $(100000,0.0001,100,83)$ $(0.1,100,0.1,203)$
parkinsons $71.19$ $71.19$ $81.36$ $77.97$ $89.83$ $77.97$ $93.1$ $87.4$ $88.21$ $89.83$ $93.22$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.1,163)$ $(1,63)$ $(0.1,203)$ $(0.1,163)$ $(0.25,0.25,1,2)$ $(10,1000,100)$ $(1,1,0.0001,0.000001)$ $(1000,100000,0.1,163)$ $(1,1000,0.01,203)$
pima $76.19$ $73.33$ $74.89$ $74.03$ $74.46$ $74.03$ $69.13$ $74.9$ $72.91$ $76.19$ $74.89$
$(0.01)$ $(0.00001,0.00001,0.00001)$ $(0.1,63)$ $(0.01,143)$ $(0.01,143)$ $(0.1,63)$ $(0.25,2,2,4)$ $(1000,10,50)$ $(10,2.5,0.0001,0.001)$ $(0.001,1,0.01,3)$ $(10000,0.01,0.001,3)$
pittsburg_bridges_T_OR_D $70.85$ $65.85$ $89.32$ $93.55$ $80.65$ $80.65$ $90$ $82.52$ $80.98$ $90.32$ $74.19$
$(100)$ $(0.00001,0.00001,0.00001)$ $(0.01,83)$ $(0.01,43)$ $(1000,3)$ $(0.01,3)$ $(0.25,0.25,1,4)$ $(10000,1,70)$ $(0.01,5,0.000001,0.00001)$ $(0.00001,1,0.01,163)$ $(10000,10,0.0001,3)$
planning $65.85$ $63.64$ $76.36$ $76.36$ $76.36$ $76.36$ $68.52$ $62.87$ $65.65$ $76.36$ $63.64$
$(10)$ $(0.00001,0.00001,0.00001)$ $(0.01,23)$ $(0.00001,3)$ $(0.00001,23)$ $(0.01,23)$ $(0.25,0.25,1,1)$ $(1,0.0001,100)$ $(2.5,10,0.000001,0.00001)$ $(0.01,0.00001,0.0001,23)$ $(0.0001,0.01,100000,43)$
ripley $89.07$ $80.67$ $86.13$ $87.53$ $87.53$ $87.53$ $89.07$ $85.17$ $86.84$ $89.87$ $89.6$
$(100000)$ $(0.00001,0.00001,0.00001)$ $(10000,103)$ $(100,203)$ $(10000,123)$ $(10000,103)$ $(0.25,0.25,0.25,4)$ $(10,1,10)$ $(5,10,0.0001,0.001)$ $(1,0.00001,0.0001,123)$ $(10000,0.01,10,83)$
Average Acc. $76.65$ $67.24$ $82.66$ $81.7$ $83.14$ $81.26$ $79.9$ $81.1$ $81.98$ $85.15$ $83.18$
Average Rank $7.72$ $10.5$ $5.54$ $6.26$ $4.94$ $6.06$ $6.52$ $6.7$ 66 $2.15$ $3.61$